How do we evaluate Sin 10 - Cos 10?


 


regards


 

Grade:12

1 Answers

ruchi yadav
10 years ago
Question should besin 20 cos 10 + cos 20 sin 10
remember sin (a+b) = sin a cos b + cos a sin b ?

sin 20 cos 10 + cos 20 sin 10
= sin (20+10)
= sin 30
= 1/2

Thank You
Ruchi
Askiitians Faculty